# Large Stepsize Gradient Descent for Logistic Loss: Non-Monotonicity of the Loss Improves Optimization Efficiency

### v1

@@ 3 SGD for Logistic Regression
Assumption 2 (Bounded and separable distribution).
@@
Theorem 4 (Risk and error bounds).
@@
The proof of Theorem 4 is deferred to Appendix E. The population logistic loss bound in Theorem 4 allows SGD with a stepsize as large as $o(t)$. For instance, in $t$ steps, SGD achieves an $\mathcal{O}(1/\sqrt{t})$ population logistic loss on average with $\eta:=\sqrt{t}$. This is in contrast to the typical online smooth convex optimization results, where the stepsize for SGD is a small constant or even vanishing (see Hazan, 2022, Chapter 3, for example).
The population zero-one error bound in Theorem 4 allows SGD with an arbitrarily large stepsize. In particular, for a large stepsize $\eta\geq\ln(\gamma^{2}t)$, there exists $\hat{\mathbf{w}}\in(\mathbf{w}_{k})_{k=0}^{t-1}$ such that
@@
Unlike the result in Section 2 for batch GD, Theorem 4 does not show the benefits of large stepsize for SGD (except on logarithmic factors). Specifically, the logistic loss bound in Theorem 4 is minimized when $\eta=\Theta\big{(}\ln(t)\big{)}$ (and larger stepsize such as $\eta\geq\omega\big{(}\ln(t)\big{)}$ hurts the bound), and the zero-one error bound in Theorem 4 is minimized when $\eta\geq\Omega\big{(}\ln(t)\big{)}$ (so larger stepsize is as good). However, both optimized bounds do not improve the bounds for $\eta=\Theta(1)$ ignoring logarithmic factors. We leave it as an open problem to investigate whether or not a large stepsize benefits SGD.
@@ 4 General Loss Functions and Nonlinear Predictors
Assumption 3 (Loss conditions).
@@
Proposition 5 provides three examples of loss functions. The proof is included in Appendix F.
Proposition 5 (Examples).
@@
where the trainable parameters are denoted by a vector $\mathbf{w}\in\mathbb{R}^{md}$, a stack of $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(m)}\in\mathbb{R}^{d}$. Here, we follow the standard NTK setup (Du et al., 2018; Ji and Telgarsky, 2019) and assume that $(a_{s})_{s=1}^{m}$ are fixed parameters satisfying [^2] Some papers assume that $(a_{s})_{s=1}^{m}$ are uniformly sampled from {±$\{\pm 1\}$. In this case, $\big{|}\sum_{s=1}^{m}a_{s}\big{|}\leq\sqrt{2m\ln(2/\delta)}$ with probability at least $1-\delta$. So we recover theirs by setting $C_{a}:=\sqrt{2\ln(2/\delta)}$ in (5) and applying a union bound.
@@
Theorem 6 (General losses and NTK).
@@
[…]
@@
[…]
@@
The proof of Theorem 6 is deferred to Appendix G. Theorem 6 characterizes the behaviors of large stepsize GD for a wide two-layer network under general classification losses. We remark that the proof of Theorem 6 directly adapts when replacing the two-layer network with a linear predictor. For completeness, we state a variant of Theorem 6 for GD in linear classification under general losses as Theorem 35 in Appendix H.
@@
The final convergence rate of GD in Theorem 6 depends on the loss functions and the stepsizes. We summarize and compare several key examples in Table 1, where we assume the total number of steps $T$ is large and treat all other instance specific quantities (such as $n$ and $\gamma$) as constants. Results in Table 1 are direct consequences of Theorem 6 and Proposition 5. In particular, we see that GD attains an improved loss by using a large stepsize that balances the length of the EoS and the stable phases.
Table 1: Effects of the loss functions and stepsizes in Theorem 6.
@@
Notably, Theorem 6 (also Theorem 35 in Appendix H) reflects the role of each loss function condition in Assumption 3. Specifically, a general classification loss is specified by 3A. Then the EoS phase bound in Theorem 6 holds when the loss satisfies an additional Lipschitz condition (3B), and the stable phase bound holds if the loss further satisfies the self-boundedness condition (3C). Finally, an exponential tail condition (3D) implies a better phase transition bound.
@@
The width condition in Theorem 6 depends on the loss function $\ell$ and the stepsize $\eta$. When we specialize $\ell$ to the logistic loss (so $\rho(\lambda)=\mathcal{O}(\ln^{2}(\lambda))$ by Proposition 5) and the stepsize to a constant ($\eta=\Theta(1)$), the width condition is polylogarithmic in the number of samples and the number of iterates, and the achieved convergence rate is $\tilde{\mathcal{O}}(1/t)$. This recovers the main result of Ji and Telgarsky (2019).
On the other hand, the width condition in Theorem 6 is at least $\Omega(\eta^{2})$, which will become polynomial in the number of steps when GD is equipped with a polynomially large stepsize $\eta=\mathrm{poly}(T)$. This is in stark contrast to the previous example where $\eta=\Theta(1)$ and the width only needs to be polylogarithmic in $T$, suggesting that a larger stepsize helps GD escape the NTK regime in the early optimization phase. We leave it as a future direction to study the early-phase feature learning of GD caused by a large stepsize.
The loss function affects the width condition in Theorem 6 through $\rho(\gamma^{2}\eta T)$. For instance, the width condition becomes $m\geq\Omega(T^{2/(a+2)})$ for the constant stepsize and the flattened polynomial loss of degree $a>0$ (see Table 1). Such results are new in the NTK literature to our knowledge.
@@ 5 Conclusion
We study constant-stepsize GD for training linear and nonlinear predictors (in the NTK regime) under general classification loss functions and constant-stepsize SGD for logistic regression, assuming suitable separability conditions on the dataset. We show GD and SGD converge even with a large stepsize that leads to a locally oscillatory loss. Moreover, we show that a large stepsize allows GD to attain an accelerated loss by undergoing an unstable initial phase, which cannot be attained if the stepsize is small such that GD converges monotonically. We leave for future work to relax the separability conditons.
Acknowledgement
We thank Fabian Pedregosa for his suggestions on an early draft. We gratefully acknowledge the support of the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639 respectively, NSF Grant DMS 2015341 and NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI).
Appendix A Proof of Theorem 1
@@
Lemma 7 (A split optimization bound).
@@
Proof of Lemma 7.
Using the definition of GD, we have
@@
which completes the proof. ∎
@@
Lemma 8 (Parameter and risk bounds in the EoS phase).
@@
Proof of Lemma 8.
In Lemma 7, choose
@@
We then apply Lemma 7 to get
@@
These complete the proof. ∎
The following lemma controls the gradient potential. We remark that the bound on gradient potential in Lemma 9 does not scale with $\eta$, while the bound on risk in Lemma 8 linearly scales with $\eta$. This difference will be crucial for getting a sharp bound on the phase transition time.
Lemma 9 (Gradient potential bound in the EoS phase).
@@
Proof of Lemma 9.
This is from the perceptron argument (Novikoff, 1962). Specifically,
@@
We complete the proof by applying the parameter bound in Lemma 8. ∎
@@
Lemma 10 (Stable phase).
@@
Proof of Lemma 10.
We verify that $\ln L(\mathbf{w})$ is $1$-smooth (this is used by Ji and Telgarsky (2018)) by direct computation:
@@
We complete the proof by induction and the condition that $L(\mathbf{w}_{s})\leq 2/\eta$. ∎
@@
Lemma 11 (A risk bound in the stable phase).
@@
[…]
Proof of Lemma 11.
The assumption enables Lemma 10 for all $t\geq s$. Therefore we have
@@
We complete the proof using the monotonicity of $L(\mathbf{w}_{t})$ for $t\geq s$ by Lemma 10. ∎
The following lemma provides a bound on the phase transition time. The bound is linear in $\eta$ because we use the gradient potential bound in Lemma 9 instead of the risk bound in Lemma 8.
Lemma 12 (Phase transition time).
@@
Proof of Lemma 12.
Applying Lemma 9 with $t=\tau$, we have
@@
which completes the proof. ∎
The above lemmas imply Theorem 1.
Proof of Theorem 1.
It follows from Lemmas 8, 11 and 12. ∎
Appendix B Proof of Corollary 2
Proof of Corollary 2.
We first verify $\tau\leq T/2$ for
@@ Appendix B Proof of Corollary 2
Next, we apply Theorem 1 with $s=\tau\leq T/2$ and get
@@
which completes the proof. ∎
Appendix C Proof of Theorem 3
@@
Lemma 13 (A lower bound).
@@
$L(\mathbf{w}_{t})\geq\frac{c_{0}}{\eta t},\quad t\geq 1,$
@@
Proof of Lemma 13.
Assume $y_{i}=1$ without loss of generality. Define
@@
$\|\bar{\mathbf{w}}_{t}\|\leq\max\{4n/b,\eta n/b\}+\eta=:c_{1}.$
By Lemma B.7 in (Wu et al., 2023), we get
@@
$G_{\max}+H_{\max}$ $=\sup_{\|\bar{\mathbf{w}}_{t}\|\leq c_{1}}\sum_{i=1}^{n}\exp(-\bar{\mathbf{x}}_{i}^{\top}\bar{\mathbf{w}})\leq n\exp(c_{1}).$
@@
$w_{t}\leq\frac{1}{\gamma}\ln\big{(}e\gamma\eta\exp(c_{1})(t+1)\big{)},\quad t\geq 0.$
@@
$\geq\ln\big{(}1+\exp(-\gamma w_{t}-c_{1})\big{)}$
$\geq\exp(-\gamma w_{t}-c_{1})$
$=\exp(-\ln\big{(}e\gamma\eta\exp(2c_{1})(t+1))\big{)}$
$\geq\frac{1}{2e\gamma\eta\exp(2c_{1})(t+1)}$
$\geq\frac{c_{0}}{\eta t},$
where $c_{0}=1/(4e\gamma\exp(2c_{1}))$ is independent of $\eta$ and $t$. ∎
@@
Proof of Lemma 14.
Since $\mathbf{w}_{0}=0,$ we have $L(0)=\ln(2)$ and
@@
This proves the original claim by contradiction. ∎
The proof of Theorem 3 is a combination of Lemmas 14 and 13.
Proof of Theorem 3.
We verify that our dataset satisfies the conditions in Lemma 14 with
@@
So by Lemma 14 and the monotonic loss, we have
@@
Then we verify that our dataset satisfies the conditions in Lemma 13 with $b=0.5.$ So by Lemma 13 and the upper bound on $\eta$, we have
$L(\mathbf{w}_{t})\geq\frac{c_{0}}{\eta t}\geq\frac{c_{1}}{t},$
where $c_{1}=c_{0}/(40\ln(2))$ is independent of $\eta$ and $t$. ∎
Appendix D A Uniform-Convergence Generalization Bound
Proposition 15 (Uniform convergence).

### v2

@@ 3 SGD for Logistic Regression
Assumption 2 (Bounded and separable distribution)
@@
Theorem 3.1 (Risk and error bounds).
@@
The proof of Theorem 3.1 is deferred to Appendix F. The population logistic loss bound in Theorem 3.1 allows SGD with a stepsize as large as $o(t)$. For instance, in $t$ steps, SGD achieves an $\mathcal{O}(1/\sqrt{t})$ population logistic loss on average with $\eta:=\sqrt{t}$. This is in contrast to the typical online smooth convex optimization results, where the stepsize for SGD is a small constant or even vanishing (see Hazan, 2022, Chapter 3, for example).
The population zero-one error bound in Theorem 3.1 allows SGD with an arbitrarily large stepsize. In particular, for a large stepsize $\eta\geq\ln(\gamma^{2}t)$, there exists $\hat{\mathbf{w}}\in(\mathbf{w}_{k})_{k=0}^{t-1}$ such that
@@
Unlike the result in Section 2 for batch GD, Theorem 3.1 does not show the benefits of a large stepsize for SGD (except on logarithmic factors). Specifically, the logistic loss bound in Theorem 3.1 is minimized when $\eta=\Theta\big{(}\ln(t)\big{)}$ (and larger stepsize such as $\eta\geq\omega\big{(}\ln(t)\big{)}$ hurts the bound), and the zero-one error bound in Theorem 3.1 is minimized when $\eta\geq\Omega\big{(}\ln(t)\big{)}$ (so larger stepsize is as good). However, both optimized bounds do not improve the bounds for $\eta=\Theta(1)$ ignoring logarithmic factors. We leave it as an open problem to investigate whether or not a large stepsize benefits SGD.
@@ 4 General Loss Functions and Nonlinear Predictors
Assumption 3 (Loss conditions)
@@
Proposition 4.1 provides three examples of loss functions. The proof is included in Appendix G.
Proposition 4.1 (Examples).
@@
where the trainable parameters are denoted by $\mathbf{w}\in\mathbb{R}^{md}$, a stack of $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(m)}\in\mathbb{R}^{d}$. We define $\phi^{\prime}(0):=0$ for concreteness but our analysis can be easily generalized to other subderivatives, that is, $\phi^{\prime}(0)=c$ for any $c\in[0,1]$. Here, we follow the standard NTK setup (Du et al., 2018; Ji and Telgarsky, 2019) and assume that $(a_{s})_{s=1}^{m}$ are fixed parameters satisfying [^2] Some papers assume that $(a_{s})_{s=1}^{m}$ are uniformly sampled from {±$\{\pm 1\}$. In this case, $\big{|}\sum_{s=1}^{m}a_{s}\big{|}\leq\sqrt{2m\ln(2/\delta)}$ with probability at least $1-\delta$. So we recover theirs by setting $C_{a}:=\sqrt{2\ln(2/\delta)}$ in (5) and applying a union bound.
@@
Theorem 4.2 (General losses and NTK).
@@
$m\geq\bigg{(}\frac{30R^{1/3}+10\ln^{1/4}(n/\delta)}{\gamma}\bigg{)}^{6},\quad\text{where}\ \ R:=6\frac{\sqrt{\rho(\gamma^{2}\eta T)}+C_{a}+\sqrt{2\ln(2n/\delta)}+\eta C_{g}}{\gamma}.$
@@
    $\tau:=\frac{1}{\gamma^{2}}\max\bigg{\{}\frac{\psi^{-1}\big{(}C_{1}(\eta+n)\big{)}}{\eta},\ C_{1}(\eta+n)\eta\bigg{\}},\ \ \text{where}\ \ \psi(\lambda):=\frac{\lambda}{\rho(\lambda)}.$
@@
The proof of Theorem 4.2 is deferred to Appendix H. Theorem 4.2 characterizes the behaviors of large stepsize GD for a wide two-layer network under general classification losses. We remark that the proof of Theorem 4.2 directly adapts when replacing the two-layer network with a linear predictor. For completeness, we state a variant of Theorem 4.2 for GD in linear classification under general losses as Theorem I.1 in Appendix I.
@@
The final convergence rate of GD in Theorem 4.2 depends on the loss functions and the stepsizes. We summarize and compare several key examples in Table 1, where we assume the total number of steps $T$ is large and treat all other instance specific quantities (such as $n$ and $\gamma$) as constants. Results in Table 1 are direct consequences of Theorem 4.2 and Proposition 4.1. In particular, we see that GD attains an improved loss by using a large stepsize that balances the length of the EoS and the stable phases.
Table 1: Effects of the loss functions and stepsizes in Theorem 4.2.
@@
Notably, Theorem 4.2 (also Theorem I.1 in Appendix I) reflects the role of each loss function condition in Assumption 3. Specifically, a general classification loss is specified by 3A. Then the EoS phase bound in Theorem 4.2 holds when the loss satisfies an additional Lipschitz condition (3B), and the stable phase bound holds if the loss further satisfies the self-boundedness condition (3C). Finally, an exponential tail condition (3D) implies a better phase transition bound.
@@
The width condition in Theorem 4.2 depends on the loss function $\ell$ and the stepsize $\eta$. When we specialize $\ell$ to the logistic loss (so $\rho(\lambda)=\mathcal{O}(\ln^{2}(\lambda))$ by Proposition 4.1) and the stepsize to a constant ($\eta=\Theta(1)$), the width condition is polylogarithmic in the number of samples and the number of iterates, and the achieved convergence rate is $\tilde{\mathcal{O}}(1/t)$. This recovers the main result of Ji and Telgarsky (2019).
On the other hand, the width condition in Theorem 4.2 is at least $\Omega(\eta^{2})$, which will become polynomial in the number of steps when GD is equipped with a polynomially large stepsize $\eta=\mathrm{poly}(T)$. This is in stark contrast to the previous example where $\eta=\Theta(1)$ and the width only needs to be polylogarithmic in $T$, suggesting that a larger stepsize helps GD escape the NTK regime in the early optimization phase. We leave it as a future direction to study the early-phase feature learning of GD caused by a large stepsize.
The loss function affects the width condition in Theorem 4.2 through $\rho(\gamma^{2}\eta T)$. For instance, the width condition becomes $m\geq\Omega(T^{2/(a+2)})$ for the constant stepsize and the flattened polynomial loss of degree $a>0$ (see Table 1). Such results are new in the NTK literature to our knowledge.
@@ 5 Conclusion
We study constant-stepsize GD for training linear and nonlinear predictors (in the NTK regime) under general classification loss functions and constant-stepsize SGD for logistic regression, assuming suitable separability conditions on the dataset. We show GD and SGD converge even with a large stepsize that leads to a locally oscillatory loss. Moreover, we show that a large stepsize allows GD to attain an accelerated loss by undergoing an unstable initial phase, which cannot be attained if the stepsize is small such that GD converges monotonically. We leave for future work to relax the separability conditions.
[…]
We thank the anonymous reviewers and area chairs for their helpful comments. We thank Fabian Pedregosa for his suggestions on an early draft. We gratefully acknowledge the support of the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639 respectively, NSF Grant DMS 2015341 and NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI).
Appendix A Additional Simulation
In Figure 2, we recreate Figure 1 with training accuracy curves. The plots show that the stable phase is entered before reaching $100\%$ training accuracy. Moreover, GD with larger stepsizes reaches perfect training accuracy faster. The simulations suggest that the phase transition time bound given in Theorem 2.1, especially the dependence on the sample size $n$, might be improvable.
Figure 2: Training loss and training accuracy of GD for logistic regression on a subset of MNIST. This is a replication of Figure 1 with training accuracy curves. The plots show that the stable phase is entered before reaching $100\%$ training accuracy. Moreover, GD with larger stepsizes reaches perfect training accuracy faster.
Appendix B Proof of Theorem 2.1
@@ Appendix A Additional Simulation
Lemma B.1 (A split optimization bound).
@@
Proof B.2.
(of Lemma B.1 Using the definition of GD, we have
@@
which completes the proof.
@@
Lemma B.3 (Parameter and risk bounds in the EoS phase).
@@
Proof B.4.
(of Lemma B.3 In Lemma B.1, choose
@@
We then apply Lemma B.1 to get
@@
These complete the proof.
The following lemma controls the gradient potential. We remark that the bound on gradient potential in Lemma B.5 does not scale with $\eta$, while the bound on risk in Lemma B.3 linearly scales with $\eta$. This difference will be crucial for getting a sharp bound on the phase transition time.
Lemma B.5 (Gradient potential bound in the EoS phase).
@@
Proof B.6.
(of Lemma B.5 This is from the perceptron argument (Novikoff, 1962). Specifically,
@@
We complete the proof by applying the parameter bound in Lemma B.3.
@@
Lemma B.7 (Stable phase).
@@
Proof B.8.
(of Lemma B.7 We verify that $\ln L(\mathbf{w})$ is $1$-smooth (this is used by Ji and Telgarsky (2018)) by direct computation:
@@
We complete the proof by induction and the condition that $L(\mathbf{w}_{s})\leq 2/\eta$.
@@
Lemma B.9 (A risk bound in the stable phase).
@@
$L(\mathbf{w}_{t})\leq\frac{2F(\mathbf{w}_{s})+\ln^{2}(\gamma^{2}\eta(t-s))}{\gamma^{2}\eta(t-s)},\ \ \text{where}\ \ F(\mathbf{w}):=\frac{1}{n}\sum_{i=1}^{n}\exp(-\mathbf{x}_{i}^{\top}\mathbf{w}).$
Proof B.10.
(of Lemma B.9 The assumption enables Lemma B.7 for all $t\geq s$. Therefore we have
@@
We complete the proof using the monotonicity of $L(\mathbf{w}_{t})$ for $t\geq s$ by Lemma B.7.
The following lemma provides a bound on the phase transition time. The bound is linear in $\eta$ because we use the gradient potential bound in Lemma B.5 instead of the risk bound in Lemma B.3.
Lemma B.11 (Phase transition time).
@@
Proof B.12.
(of Lemma B.11 Applying Lemma B.5 with $t=\tau$, we have
@@
which completes the proof.
The above lemmas imply Theorem 2.1.
Proof B.13.
(of Theorem 2.1 It follows from Lemmas B.3, B.9 and B.11.
Appendix C Proof of Corollary 2.2
Proof C.1.
(of Corollary 2.2 We first verify $\tau\leq T/2$ for
@@
Next, we apply Theorem 2.1 with $s=\tau\leq T/2$ and get
@@
which completes the proof.
Appendix D Proof of Theorem 2.3
@@
Lemma D.1 (A lower bound).
@@
$L(\mathbf{w}_{t})\geq\frac{1}{c_{0}\eta\exp(c_{0}\eta)t},\quad t\geq 1,$
@@
Proof D.2.
(of Lemma D.1 Assume $y_{i}=1$ without loss of generality. Define
@@
$\|\bar{\mathbf{w}}_{t}\|\leq\max\{4n/b,\eta n/b\}+\eta\leq c_{1}\eta,$
where $c_{1}$ is a positive constant independent of $\eta$. By Lemma B.7 in (Wu et al., 2023), we get
@@
$G_{\max}+H_{\max}$ $=\sup_{\|\bar{\mathbf{w}}_{t}\|\leq c_{1}\eta}\sum_{i=1}^{n}\exp(-\bar{\mathbf{x}}_{i}^{\top}\bar{\mathbf{w}})\leq n\exp(c_{1}\eta).$
@@
$w_{t}\leq\frac{1}{\gamma}\ln\big{(}e\gamma\eta\exp(c_{1}\eta)(t+1)\big{)},\quad t\geq 0.$
@@
$\geq\ln\big{(}1+\exp(-\gamma w_{t}-c_{1}\eta)\big{)}$
$\geq\exp(-\gamma w_{t}-c_{1}\eta)$
$=\exp(-\ln\big{(}e\gamma\eta\exp(2c_{1}\eta)(t+1))\big{)}$
$\geq\frac{1}{2e\gamma\eta\exp(2c_{1}\eta)(t+1)}$
$\geq\frac{1}{c_{0}\eta\exp(c_{0}\eta)t},$
where $c_{0}>0$ is independent of $\eta$ and $t$.
@@
Proof D.4.
(of Lemma D.3 Since $\mathbf{w}_{0}=0,$ we have $L(0)=\ln(2)$ and
@@
This proves the original claim by contradiction.
The proof of Theorem 2.3 is a combination of Lemmas D.3 and D.1.
Proof D.5.
(of Theorem 2.3 We verify that our dataset satisfies the conditions in Lemma D.3 with
@@
So by Lemma D.3 and the monotonic loss, we have
@@
Then we verify that our dataset satisfies the conditions in Lemma D.1 with $b=0.5.$ So by Lemma D.1 and the upper bound on $\eta$, we have
$L(\mathbf{w}_{t})\geq\frac{1}{c_{0}\eta\exp(c_{0}\eta)t}\geq\frac{c_{1}}{t},$
where $c_{1}>0$ is independent of $\eta$ and $t$ since $\eta$ is upper bounded and $c_{0}>0$ is independent of $\eta$ and $t$.
Appendix E A Uniform-Convergence Generalization Bound
Proposition E.1 (Uniform convergence).